The left end of a taut string of length L, is connected to a vibrator with a fixed frequency f. The right end of the string is tied to a suspended object of varying mass, m, through a pulley. For a mass m1 of the object, a standing wave with one loop is observed. For a mass m2 of the object, a standing wave with two loops is observed. What is the ratio "m2 / m1 = ?"


0.0625


0.25


2


4


16


1
Expert's answer
2021-05-26T21:35:16-0400

Gives string length =L

Frequency=n


Mass (m)=m1

Loop(n)=1

Mass(m)=m2

Loop(n)=2

Case (1)

"L=\frac{n\lambda}{2}"

n=1

"L=\frac{\lambda}{2}"

"\lambda=2L"

"v=\sqrt\frac{T}{m_1}"

"\frac{T}{m_1}=({n\lambda})^2"

"\frac{T}{m_1}=({n\times 2L})^2"

"m_1=\frac{T}{v^2}"

"m_1=\frac{T}{4(n\lambda)^2}"

Case(2)

n=2

"L=\frac{2\lambda}{2}"

"L=\lambda"

"\frac{T}{m_2}=({n\times L})^2"

"m_2=\frac{T}{(n\lambda)^2}"

"\frac{m_2}{m_1}=\frac{\frac{T}{(n\lambda)^2}}{\frac{T}{4(n\lambda)^2}}=\frac{4}{1}"